OpenStudy (anonymous):

what does x equal, when 0=2e^(x/2)?

OpenStudy (rizags):

so, this is what you need to solve? \[0=2e^{\frac{x}{2}}\]

OpenStudy (rizags):

is this correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

without graphing calculator :)

OpenStudy (rizags):

well, begin by dividing both sides by 2 to get \[0=e^{\frac{x}{2}}\]

OpenStudy (freckles):

hint: can anything positive ever by 0?

OpenStudy (anonymous):

mhm I got that

OpenStudy (anonymous):

? @freckles huh?

OpenStudy (anonymous):

@Rizags what is next?

OpenStudy (anonymous):

i got to ln0 = x/2

OpenStudy (rizags):

which is impossible

OpenStudy (rizags):

there are no solutions in the reals

OpenStudy (freckles):

e^(raise to any power) is positive

OpenStudy (anonymous):

but I don't get what ln0 means and i typed it into the calculator, which said Error

OpenStudy (freckles):

a positive number cannot be zero

OpenStudy (anonymous):

oh what...

OpenStudy (rizags):

right, because the answer is imaginary

OpenStudy (anonymous):

but my calculus question needs the bounds to be calculated

OpenStudy (freckles):

what is the question exactly?

OpenStudy (rizags):

no solutions exist for x actually, not even imaginary ones

OpenStudy (anonymous):

That is the question

OpenStudy (freckles):

 bounds are given